Two vessels of different materials are similar in size in every respect. The same quantity of ice filled in them gets melted in 20 minutes and 40 minutes respectively. The ratio of thermal conductivities of the materials is

Grade:11

1 Answers

Arun
6 years ago
Q = K* A (\theta1 - \theta2)* t/L
Because A, L and temp. Difference
Hence
At constant pressure
K1/K2 = t1/t2
K1/K2 = 40/20 = 2